OpenStudy (anonymous):

Solve. 3x^2 = –15x + 21 -5+_ i sqrt 3 / 2 5 +_ sqrt 53 / 2 5 +_ 3 sqrt 3 / 2 -5 +_ sqrt 23 / 2

OpenStudy (mimi_x3):

use the quadratic formula x = 1.14 and x = -6.1400

OpenStudy (anonymous):

\[3x ^{2}+15x-21=0\] cancelling 3 \[x ^{2}+5x-7=0\] \[x=(-5\pm \sqrt{5^{2}-4*1*(-7)})/2\] \[-5\pm \sqrt{53}/2\]